1.ABC industries produces 1 table and sells it at shs 100. The marginal cost of production is shs. 60 and the fixed cost is shs. 4,000 per annum.

Calculate :

a) P/V ratio                                                   ( 1 marks)

b) break-even sales                                    ( 2 marks)

c) the sales to earn a profit of shs. 5,000 ( 3 marks)

d) profit at sales of shs 30,000                   ( 3 marks)


1
Expert's answer
2021-09-26T20:18:15-0400

Solution:

a.). P/V ratio = "\frac{Contribution}{Sales} \times 100"

Contribution = Sales – Variable costs = 100 – 60 = 40

P/V ratio = "\frac{40}{100} \times 100 = 40\%"

 

P/V ratio = "= 40\%"


b.). Break-even sales = "\frac{Fixed \; costs \times Sales }{Contribution}"


= "\frac{4,000 \times 100 }{40} = 10,000"

Break-even sales = shs. 10,000

 

c.). The sales to earn a profit of shs. 5,000:

Desired sales = "\frac{Fixed \; costs + Profit }{P/V ratio}"


= "\frac{4,000 + 5,000 }{0.4} = \frac{9,000}{0.4} = 22,500"

The sales to earn a profit of shs. 5,000 = shs.22,500

 

d.). Profit at sales of shs.30,000:

Sales = Fixed Costs + Profit/P/V ratio

30,000 = 4,000 + "\frac{P }{0.4}"

30,000 "\times" 0.4 = 4,000 + P

12,000 = 4,000 + P

P = 12,000 – 4,000 = 8,000

Profit at sales of shs.30,000 = shs.8,000